# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ Appendix B Canonical 2-Edge Cover
This section is dedicated to the proof of Definition 7, which we first restate.
@@
We first compute a triangle-free $2$-edge cover $H^{\prime}$ (which can be done in polynomial time due to Lemma 4) and show how to modify it to obtain a canonical $2$-edge cover with the same number of edges. Since $H^{\prime}$ is a triangle-free $2$-edge cover, each component is either a 2EC component that contains at least $4$ edges, or complex.
If $H^{\prime}$ is canonical, we are done. Otherwise, we apply the following algorithm to $H^{\prime}$: While there exist sets of edges $F_{A}\subseteq E\setminus H^{\prime}$ and $F_{R}\subseteq H^{\prime}$ with $|F_{A}|\leq|F_{R}|\leq 2$ such that $H^{\prime\prime}\coloneq(H^{\prime}\setminus F_{R})\cup F_{A}$ is a triangle-free $2$-edge cover and either $|H^{\prime\prime}|<|H^{\prime}|$, or $|H^{\prime\prime}|=|H^{\prime}|$ and $H^{\prime\prime}$ contains strictly less connected components than $H^{\prime}$, or $H^{\prime\prime}$ contains strictly less bridges than $H^{\prime}$, or $H^{\prime\prime}$ contains less cut vertices in 2EC components than $H^{\prime}$, we update $H^{\prime}\coloneq H^{\prime\prime}$.
Clearly, each iteration of this algorithm can be executed in polynomial time. Moreover, there can only be polynomially many iterations. Thus, let $H$ be equal to $H^{\prime}$ when the algorithm terminates. We prove that $H$ is a canonical 2-edge cover of $G$. This proves the lemma because the algorithm does not increase the number of edge compared to a minimum triangle-free $2$-edge cover of $G$. To this end, we consider the following cases.
@@
Case 1.1: $C$ has a cut vertex $v$. Observe that $v$ must have degree $4$ in $C$. Furthermore, each connected component in $C\setminus\{v\}$ contains at least $2$ vertices, as otherwise $C$ contains parallel edges, a contradiction to $G$ being structured. There must be a connected component $K_{1}$ in $C\setminus\{v\}$ such that $K_{1}$ contains at most 2 vertices (and hence at most 1 edge). This is true as otherwise each connected component has at least 2 edges in $C\setminus\{v\}$ and since there are at least $2$ connected components in $C\setminus\{v\}$, because $v$ is a cut vertex, $C$ must contain at least $8$ edges, a contradiction. Assume that there is a component $K_{1}$ with $2$ vertices and let $u_{1}$ and $u_{2}$ be these vertices. Note that the three edges $\{vu_{1},u_{1}u_{2},u_{2}v\}$ must be in $C$ as otherwise $G$ is not structured since it must contain parallel edges. If there exists an edge $u_{1}w\in E\setminus H$ (or $u_{2}w\in E$ by symmetry) for $w\in V\setminus V(C)$, then we can remove $vu_{1}$ from $H$ and add $u_{1}w$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$ with strictly less components and $|H|=|H^{\prime\prime}|$, a contradiction to the algorithm. If there exists an edge $u_{1}w\in E\setminus H$ (or $u_{2}w\in E$ by symmetry) for $w\in V(C)\setminus\{v,u_{1},u_{2}\}$, then we can remove $vu_{1}$ from $H$ and add $u_{1}w$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$ with $|H^{\prime\prime}|=|H|$ in which $v$ is not a cut vertex anymore. If neither of these two cases apply, we conclude $\deg_{G}(u_{i})=2$ for $i=1,2$. Thus, $v-u_{1}-u_{2}-v$ is a contractible cycle, a contradiction to $G$ being structured.
@@
Case 1.2: $C$ contains no 2-vertex cut (and, thus, $C$ is 3-vertex-connected ). This implies that every vertex has degree at least 3 in $C$, as otherwise the neighboring vertices of a vertex of $C$ with degree $2$ form a $2$-vertex cut. But then it must hold that $|V(C)|=4$, since $|V(C)|\leq 3$ is not possible with the above degree constraint and if $|V(C)|\geq 5$, then the number of edges in $C$ is at least $8$, a contradiction. But if $|V(C)|=4$ and $C$ is $3$-vertex-connected, $C$ must be a complete graph on $4$ vertices. Therefore, there is an edge $e\in E(C)$ such that $H\setminus\{e\}$ is a triangle-free 2-edge cover, a contradiction to the algorithm.
@@
Case 1.3: $C$ contains a 2-vertex cut $\{v_{1},v_{2}\}$. Note that each connected component of $C\setminus\{v_{1},v_{2}\}$ must have an edge to both $v_{1}$ and $v_{2}$, as otherwise $C$ contains a cut vertex and we are in Case 1.1. Hence, there are at most $3$ connected components in $C\setminus\{v_{1},v_{2}\}$, as otherwise the number of edges in $C$ is at least 8.
@@
First, assume that $C\setminus\{v_{1},v_{2}\}$ contains exactly $3$ connected components. Then there can be at most one connected component in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices, as otherwise the total number of edges in $C$ exceeds $7$. In any case, each vertex $u$ of some connected component in $C\setminus\{v_{1},v_{2}\}$ must have an edge $e_{u}$ to $\{v_{1},v_{2}\}$ as otherwise $C$ contains a cut vertex and we are actually in Case 1.1. If there is a vertex $u$ of some connected component in $C\setminus\{v_{1},v_{2}\}$ that has an edge $uw$ in $G$ to some vertex $w\in V\setminus V(C)$, then we can replace $e_{u}$ with $uw$ in $H$ and thereby decrease the number of connected components, a contradiction to the algorithm. Hence, assume that this is not true. Assume that there is exactly one connected component $K_{1}$ in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices $x_{1}$ and $x_{2}$. Note that $K_{1}$ must consist of the single edge $x_{1}x_{2}$ as otherwise we have at least $8$ edges in $C$. Let $K_{2}$ and $K_{3}$ be the other two connected components in $C\setminus\{v_{1},v_{2}\}$ consisting of single vertices $y$ and $z$, respectively. By the above arguments, we have that $C$ contains the edges $\{v_{1}x_{1},x_{1}x_{2},x_{2}v_{2},v_{1}y,yv_{2},v_{1}z,zv_{2}\}$ (up to symmetry). If there is an edge $x_{1}y$ in $G$ (or any of $\{x_{1}z,x_{2}y,x_{2}z,yz\}$ by symmetry), then we can remove $v_{2}y$ and $x_{1}v_{1}$ from $H$ and add $x_{1}y$ to obtain a triangle-free $2$-edge cover with strictly less edges than $H$, a contradiction to the algorithm. Hence, assume that none of the edges exists in $G$. But then all vertices in $\{x_{1},x_{2},y,z\}$ have degree $2$ in $G$ and, thus, $C$ is contractible, a contradiction. The case that there is no component in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices, i.e., in which each component consists of exactly a single vertex, is similar. This finishes the case that in $C\setminus\{v_{1},v_{2}\}$ there are exactly 3 connected components.
@@
Hence, assume that $C\setminus\{v_{1},v_{2}\}$ contains exactly 2 connected components $K_{1}$ and $K_{2}$ and without loss of generality assume $|E(K_{1})|\geq|E(K_{2})|$. Recall that each connected component of $C\setminus\{v_{1},v_{2}\}$ must have an edge to both $v_{1}$ and $v_{2}$. Hence, $\{v_{1},v_{2}\}$ is incident to at least 4 edges in $C$. Thus, $|E(K_{2})|\leq|E(K_{1})|\leq 3$. If $|E(K_{1})|=3$, then $K_{1}$ is a path on 4 vertices. Otherwise, $K_{1}$ is a triangle on 3 vertices or a star on 4 vertices (one center and 3 leaves). If $K_{1}$ is a triangle, there must be an edge $e$ in $C$ such that $H\setminus\{e\}$ is a triangle-free $2$-edge cover, a contradiction to the algorithm. If $K_{1}$ is a star on 4 vertices, then there must be 3 edges from $K_{1}$ to $\{v_{1},v_{2}\}$, as otherwise $H$ is not a 2-edge cover. But then $C$ contains at least 8 edges, a contradiction. Hence, if $|E(K_{1})|=3$, then $K_{1}$ is a path on 4 vertices. But then one leaf of $K_{1}$ must be adjacent to $v_{1}$ in $C$ and the other leaf of $K_{1}$ must be adjacent to $v_{2}$ in $C$. Since $K_{2}$ is a single vertex in this case and must be incident to both $v_{1}$ and $v_{2}$, $C$ is a cycle on $7$ vertices, a contradiction. Hence, we have that $|E(K_{2})|\leq|E(K_{1})|\leq 2$. Since $G$ does not contain parallel edges, $K_{1}$ must be a simple path on $i$ vertices for some $i\in\{1,2,3\}$. The same is true for $K_{2}$. By similar reasoning as before we can observe that the leaves of $K_{1}$ and $K_{2}$, respectively, must be incident to $v_{1}$ and $v_{2}$ (or, if $K_{1}$ or $K_{2}$ is a single vertex, then it must be incident to both $v_{1}$ and $v_{2}$) and therefore we observe that $C$ is a simple cycle, a contradiction.
@@
Case 2: There is some complex component $C$ of $H$ containing a pendant block $B$ with less than 6 edges. Assume there is a pendant block $B$ with exactly 5 edges. The cases in which $B$ has 3 or 4 edges is similar. Note that $B$ must contain either 4 or 5 vertices. If $B$ contains exactly $4$ vertices then $B$ must contain a simple cycle on 4 vertices with an additional edge $e$ being a chord of this cycle, which is redundant, and, thus, a contradiction to the algorithm. Hence, assume that $B$ contains exactly $5$ vertices, i.e., $B$ is a cycle $b_{1}-b_{2}-b_{3}-b_{4}-b_{5}-b_{1}$, and assume that $b_{1}$ is incident to the unique bridge $f$ in $C$. If either $b_{2}$ (or $b_{5}$ by symmetry) is incident to an edge $e$ with neighbor in $V\setminus V(B)$ we can remove $b_{1}b_{2}$ from $H$ and add $e$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$. If $e$ is incident to some vertex in $V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $e$ is incident to some vertex in $V(C)\setminus V(B)$, or $H^{\prime\prime}$ has strictly less bridges compared to $H$. This is a contradiction to the algorithm. Thus, in the following we can assume that $b_{2}$ and $b_{5}$ are only adjacent to vertices of $V(B)$. Since $G$ is structured, $b_{3}$ and $b_{4}$ both must be incident to some edge $e_{3}$ and $e_{4}$ going to some vertex in $V\setminus V(B)$, respectively, as otherwise $B$ is contractible. Moreover, there must be the edge $b_{2}b_{5}$ in $G$, as otherwise $B$ is contractible. But then we can remove $b_{2}b_{3}$ and $b_{1}b_{5}$ from $H$ and add $b_{2}b_{5}$ and $e_{3}$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$. If $e_{3}$ is incident to some vertex $V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $e_{3}$ is incident to some vertex $V(C)\setminus V(B)$, $H^{\prime\prime}$ has strictly less bridges compared to $H$. This is a contradiction to the algorithm.
@@
Case 3: There is some complex component $C$ of $H$ containing a non-pendant block $B$ that contains less than $4$ edges, that is, $B$ is a triangle. In this case, we observe that in any case there must be an edge $e$ in $B$ such that $H\setminus\{e\}$ is a triangle-free $2$-edge cover, a contradiction to the algorithm. ∎

### v2

@@ Appendix B Canonical 2-Edge Cover
This section is dedicated to the proof of Definition˜7, which we first restate.
@@
We first compute a triangle-free $2$-edge cover $H^{\prime}$ (which can be done in polynomial time due to Lemma˜4) and show how to modify it to obtain a canonical $2$-edge cover with the same number of edges. Since $H^{\prime}$ is a triangle-free $2$-edge cover, each component is either a 2EC component that contains at least $4$ edges, or complex.
If $H^{\prime}$ is canonical, we are done. Otherwise, we apply the following algorithm to $H^{\prime}$: While there exist sets of edges $F_{A}\subseteq E\setminus H^{\prime}$ and $F_{R}\subseteq H^{\prime}$ with $|F_{A}|\leq|F_{R}|\leq 2$ such that $H^{\prime\prime}\coloneq(H^{\prime}\setminus F_{R})\cup F_{A}$ is a triangle-free $2$-edge cover and either
  • $|H^{\prime\prime}|<|H^{\prime}|$, or
  • $|H^{\prime\prime}|=|H^{\prime}|$ and either
    • $H^{\prime\prime}$ contains strictly fewer connected components than $H^{\prime}$, or
    • the number of components of $H^{\prime\prime}$ is equal to the number of components in $H^{\prime}$ and $H^{\prime\prime}$ contains strictly fewer bridges than $H^{\prime}$, or
    • the number of components of $H^{\prime\prime}$ is equal to the number of components in $H^{\prime}$, the number of bridges in $H^{\prime\prime}$ is equal to the number of bridges in $H^{\prime}$, and $H^{\prime\prime}$ contains strictly fewer cut vertices in 2EC components than $H^{\prime}$,
we update $H^{\prime}\coloneq H^{\prime\prime}$.
Clearly, each iteration of this algorithm can be executed in polynomial time. Moreover, there can only be polynomially many iterations. Thus, let $H$ be equal to $H^{\prime}$ when the algorithm terminates. We prove that $H$ is a canonical 2-edge cover of $G$. This proves the lemma because the algorithm does not increase the number of edges compared to a minimum triangle-free $2$-edge cover of $G$. To this end, we consider the following cases.
@@
Case 1.1: $C$ has a cut vertex $v$. Observe that $v$ must have degree $4$ in $C$. Furthermore, each connected component in $C\setminus\{v\}$ contains at least $2$ vertices, as otherwise $C$ contains parallel edges, a contradiction to $G$ being structured. There must be a connected component $K_{1}$ in $C\setminus\{v\}$ such that $K_{1}$ contains at most 2 vertices (and hence at most 1 edge). This is true as otherwise each connected component has at least 2 edges in $C\setminus\{v\}$ and since there are at least $2$ connected components in $C\setminus\{v\}$, because $v$ is a cut vertex, $C$ must contain at least $8$ edges, a contradiction. Assume that there is a component $K_{1}$ with $2$ vertices and let $u_{1}$ and $u_{2}$ be these vertices. Note that the three edges $\{vu_{1},u_{1}u_{2},u_{2}v\}$ must be in $C$ as otherwise $G$ is not structured since it must contain parallel edges. If there exists an edge $u_{1}w\in E\setminus H$ (or $u_{2}w\in E$ by symmetry) for $w\in V\setminus V(C)$, then we can remove $vu_{1}$ from $H$ and add $u_{1}w$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$ with strictly fewer components and $|H|=|H^{\prime\prime}|$, a contradiction to the assumption that the algorithm terminated. If there exists an edge $u_{1}w\in E\setminus H$ (or $u_{2}w\in E$ by symmetry) for $w\in V(C)\setminus\{v,u_{1},u_{2}\}$, then we can remove $vu_{1}$ from $H$ and add $u_{1}w$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$ with $|H^{\prime\prime}|=|H|$ in which $v$ is not a cut vertex anymore. Since we have not created a new cut vertex (and the number of components and number of bridges stays the same), this contradicts the assumption that the algorithm terminated. If neither of these two cases apply, we conclude $\deg_{G}(u_{i})=2$ for $i=1,2$. Thus, $v-u_{1}-u_{2}-v$ is a contractible cycle, a contradiction to $G$ being structured.
@@
Case 1.2: $C$ contains no 2-vertex cut (and, thus, $C$ is 3-vertex-connected). This implies that every vertex has degree at least 3 in $C$, as otherwise the neighboring vertices of a vertex of $C$ with degree $2$ form a $2$-vertex cut. But then it must hold that $|V(C)|=4$, since $|V(C)|\leq 3$ is not possible with the above degree constraint and if $|V(C)|\geq 5$, then the number of edges in $C$ is at least $8$, a contradiction. But if $|V(C)|=4$ and $C$ is $3$-vertex-connected, $C$ must be a complete graph on $4$ vertices. Therefore, there is an edge $e\in E(C)$ such that $H\setminus\{e\}$ is a triangle-free 2-edge cover, a contradiction to the assumption that the algorithm terminated.
@@
Case 1.3: $C$ contains a 2-vertex cut $\{v_{1},v_{2}\}$. Note that each connected component of $C\setminus\{v_{1},v_{2}\}$ must have an edge to both $v_{1}$ and $v_{2}$, as otherwise $C$ contains a cut vertex, and we are in Case 1.1. Hence, there are at most $3$ connected components in $C\setminus\{v_{1},v_{2}\}$, as otherwise the number of edges in $C$ is at least 8.
@@
First, assume that $C\setminus\{v_{1},v_{2}\}$ contains exactly $3$ connected components. Then there can be at most one connected component in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices, as otherwise the total number of edges in $C$ exceeds $7$. In any case, each vertex $u$ of some connected component in $C\setminus\{v_{1},v_{2}\}$ must have an edge $e_{u}$ to $\{v_{1},v_{2}\}$ as otherwise $C$ contains a cut vertex, and we are actually in Case 1.1. If there is a vertex $u$ of some connected component in $C\setminus\{v_{1},v_{2}\}$ that has an edge $uw$ in $G$ to some vertex $w\in V\setminus V(C)$, then we can replace $e_{u}$ with $uw$ in $H$ and thereby decrease the number of connected components, a contradiction to the algorithm. Hence, assume that this is not true. Assume that there is exactly one connected component $K_{1}$ in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices $x_{1}$ and $x_{2}$. Note that $K_{1}$ must consist of the single edge $x_{1}x_{2}$ as otherwise we have at least $8$ edges in $C$. Let $K_{2}$ and $K_{3}$ be the other two connected components in $C\setminus\{v_{1},v_{2}\}$ consisting of single vertices $y$ and $z$, respectively. By the above arguments, we have that $C$ contains the edges $\{v_{1}x_{1},x_{1}x_{2},x_{2}v_{2},v_{1}y,yv_{2},v_{1}z,zv_{2}\}$ (up to symmetry). If there is an edge $x_{1}y$ in $G$ (or any of $\{x_{1}z,x_{2}y,x_{2}z,yz\}$ by symmetry), then we can remove $v_{2}y$ and $x_{1}v_{1}$ from $H$ and add $x_{1}y$ to obtain a triangle-free $2$-edge cover with strictly fewer edges than $H$, a contradiction to the algorithm. Hence, assume that none of the edges exists in $G$. Now the vertices in $\{y,z\}$ have degree $2$ in $G$ and the vertices in $\{x_{1},x_{2}\}$ are only adjacent to $\{v_{1},v_{2},x_{1},x_{2}\}$. But then $C$ is contractible, a contradiction to $G$ being structured. The case that there is no component in $C\setminus\{v_{1},v_{2}\}$ containing 2 vertices, i.e., in which each component consists of exactly a single vertex, is similar. This finishes the case that in $C\setminus\{v_{1},v_{2}\}$ there are exactly 3 connected components.
@@
Hence, assume that $C\setminus\{v_{1},v_{2}\}$ contains exactly 2 connected components $K_{1}$ and $K_{2}$ and without loss of generality assume $|E(K_{1})|\geq|E(K_{2})|$. Recall that each connected component of $C\setminus\{v_{1},v_{2}\}$ must have an edge to both $v_{1}$ and $v_{2}$. Hence, $\{v_{1},v_{2}\}$ is incident to at least 4 edges in $C$. Thus, $|E(K_{2})|\leq|E(K_{1})|\leq 3$. If $|E(K_{1})|=3$, we prove that $K_{1}$ is a path on 4 vertices. If $K_{1}$ is not a path on $4$ vertices, then $K_{1}$ is a triangle on 3 vertices or a star on 4 vertices (one center and 3 leaves). If $K_{1}$ is a triangle, there must be an edge $e$ in $C$ such that $H\setminus\{e\}$ is a triangle-free $2$-edge cover, a contradiction to the algorithm. If $K_{1}$ is a star on 4 vertices, then there must be 3 edges from $K_{1}$ to $\{v_{1},v_{2}\}$, as otherwise $H$ is not a 2-edge cover. But then $C$ contains at least 8 edges, a contradiction. Hence, if $|E(K_{1})|=3$, then $K_{1}$ is a path on 4 vertices. But then one leaf of $K_{1}$ must be adjacent to $v_{1}$ in $C$ and the other leaf of $K_{1}$ must be adjacent to $v_{2}$ in $C$ (otherwise, there is a cut vertex and we are in Case 1.1). Since $K_{2}$ is a single vertex in this case and must be incident to both $v_{1}$ and $v_{2}$, $C$ is a cycle on $7$ vertices, a contradiction. Hence, we have that $|E(K_{2})|\leq|E(K_{1})|\leq 2$. Since $G$ does not contain parallel edges, $K_{1}$ must be a simple path on $i$ vertices for some $i\in\{1,2,3\}$. The same is true for $K_{2}$. By similar reasoning as before we can observe that the leaves of $K_{1}$ and $K_{2}$, respectively, must be incident to $v_{1}$ and $v_{2}$ (or, if $K_{1}$ or $K_{2}$ is a single vertex, then it must be incident to both $v_{1}$ and $v_{2}$) and therefore we observe that $C$ is a simple cycle, a contradiction.
@@
Case 2: There is some complex component $C$ of $H$ containing a pendant block $B$ with less than 6 edges. First, assume there is a pendant block $B$ with exactly 3 edges. In this case, $B$ is a cycle $b_{1}-b_{2}-b_{3}-b_{1}$ (since there are no parallel edges as $G$ is structured), and assume that $b_{1}$ is incident to the unique bridge $f$ in $C$. However, since $B$ is a triangle and $G$ is structured, there must be an edge $e=b_{2}w$ incident to $b_{2}$ in $G$ with $w\notin\{b_{1},b_{3}\}$ (by invoking the 3-matching Lemma˜2 to $V(B)$ and $V\setminus V(B)$). But then we can remove $b_{1}b_{2}$ and add $e$ to $H$ to obtain $H^{\prime\prime}$ and observe that $H^{\prime\prime}$ is a triangle-free 2-edge cover. If $w\in V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $w\in V(C)$, $H^{\prime\prime}$ has strictly fewer bridges compared to $H$ (and the number of connected component stays the same). This is a contradiction to the algorithm.
Next, assume there is a pendant block $B$ with exactly 4 edges. In this case, $B$ is a cycle $b_{1}-b_{2}-b_{3}-b_{4}-b_{1}$ (since there are no parallel edges as $G$ is structured), and assume that $b_{1}$ is incident to the unique bridge $f$ in $C$. If $b_{2}$ or $b_{4}$ are incident to an edge $e$ incident to some $w\in V\setminus V(B)$, then we can remove either $b_{1}b_{2}$ or $b_{1}b_{4}$ and add $e$ to $H$ to obtain $H^{\prime\prime}$ and observe that $H^{\prime\prime}$ is a triangle-free 2-edge cover. If $w\in V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $w\in V(C)$, $H^{\prime\prime}$ has strictly fewer bridges compared to $H$ (and the number of connected component stays the same). This is a contradiction to the algorithm. Hence, this is not the case. But then $\{b_{1},b_{3}\}$ is a non-isolating 2-vertex, a contradiction to $G$ being structured.
Finally, assume there is a pendant block $B$ with exactly 5 edges. Note that $B$ must contain either 4 or 5 vertices. If $B$ contains exactly $4$ vertices then $B$ must contain a simple cycle on 4 vertices with an additional edge $e$ being a chord of this cycle, which is redundant, and, thus, a contradiction to the algorithm. Hence, assume that $B$ contains exactly $5$ vertices, i.e., $B$ is a cycle $b_{1}-b_{2}-b_{3}-b_{4}-b_{5}-b_{1}$, and assume that $b_{1}$ is incident to the unique bridge $f$ in $C$. If either $b_{2}$ (or $b_{5}$ by symmetry) is incident to an edge $e$ with neighbor in $V\setminus V(B)$ we can remove $b_{1}b_{2}$ from $H$ and add $e$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$. If $e$ is incident to some vertex in $V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $e$ is incident to some vertex in $V(C)\setminus V(B)$, $H^{\prime\prime}$ has strictly fewer bridges compared to $H$ (and the number of connected component stays the same). This is a contradiction to the algorithm. Thus, in the following we can assume that $b_{2}$ and $b_{5}$ are only adjacent to vertices of $V(B)$. Since $G$ is structured, $b_{3}$ and $b_{4}$ both must be incident to some edge $e_{3}$ and $e_{4}$ going to some vertex in $V\setminus V(B)$, respectively, as otherwise $B$ is contractible. Moreover, there must be the edge $b_{2}b_{5}$ in $G$, as otherwise $B$ is contractible. But then we can remove $b_{2}b_{3}$ and $b_{1}b_{5}$ from $H$ and add $b_{2}b_{5}$ and $e_{3}$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$. If $e_{3}$ is incident to some vertex $V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $e_{3}$ is incident to some vertex $V(C)\setminus V(B)$, $H^{\prime\prime}$ has strictly fewer bridges compared to $H$ (and the number of connected component stays the same). This is a contradiction to the algorithm.
@@
Case 3: There is some complex component $C$ of $H$ containing a non-pendant block $B$ that contains less than $4$ edges, that is, $B$ is a triangle on the vertices $b_{1}-b_{2}-b_{3}$ (otherwise $B$ contains parallel edges, a contradiction to $G$ being structured). In this case, observe that either there must be an edge $e$ in $B$ such that $H\setminus\{e\}$ is a triangle-free $2$-edge cover, a contradiction to the algorithm, or all edges in $H$ incident to $V(B)$ and $V(G)\setminus V(B)$ are incident to one vertex, say $b_{1}$. But then, applying Lemma˜2, we know that there must be an edge $b_{2}w\in E(G)\setminus E(H)$, where $w\in V(G)\setminus V(B)$. But then we can remove $b_{1}b_{2}$ from $H$ and add $b_{2}w$ to obtain a triangle-free $2$-edge cover $H^{\prime\prime}$. If $w\in V\setminus V(C)$, the number of connected components in $H^{\prime\prime}$ is strictly less compared to $H$, and otherwise, that is, $w\ inV(C)\setminus V(B)$, $H^{\prime\prime}$ has strictly fewer bridges compared to $H$ (and the number of connected component stays the same). This is a contradiction to the algorithm.
Finally, observe that if $B$ contains $4$ edges but less than $4$ vertices, then there must be some parallel edge in $B$, a contradiction to the fact that $G$ is structured. ∎